OpenStudy (anonymous):

delia read 17 pages of a 68-page short story in 25 minutes . she wants to find out how long it would take her to read the entire story A)write a proportion that delia could use to find out how long it would take her to read the entire story if she continues to read at the same pace b) solve the proportion you wrote in part a . show ur work

OpenStudy (mathstudent55):

a) \[\frac{ pages }{ minutes } = \frac{ pages }{ minutes }\] \[\frac{ 17 }{ 25 } = \frac{ 68 }{ m }\] b) Cross multiply: 17m = 25(68) 17m = 1700 m = 100 Answer: 100 mintes

OpenStudy (anonymous):

thx